OpenStudy (anonymous):

What is the simplified form of (1/y)+(1/x)/(1/y)-(1/x)

OpenStudy (anonymous):

(x/xy+y/xy)/(x/xy-y/xy) =((x+y)/xy) / ((x-y) / xy) = (x+y)/(x-y)

OpenStudy (anonymous):

\[ \frac{x+y}{xy} \div \frac{x - y}{xy} = \frac{x+y}{x-y}\]

myininaya (myininaya):

\[\frac{\frac{1}{y}+\frac{1}{x}}{\frac{1}{y}-\frac{1}{x}} \cdot \frac{yx}{yx}=\frac{\frac{yx}{y}+\frac{yx}{x}}{\frac{yx}{y}-\frac{yx}{x}}\] \[=\frac{x+y}{x-y}\]